10- Show work, find the following for the function F(x)= (x+6)²(x+1)²
The x intercept
The y intercept
Does the graph cross or touch the x-axis at the smaller x intercept?
a- The graph touches x-axis
b- The graph crossed x-axis
Does the graph cross or touch the x- axis at the larger x intercept
a- Graph crosses x-axis
b- Graph touches x-axis

Answer by lwsshak3(11628)

10- Show work, find the following for the function F(x)= (x+6)²(x+1)²
F(x)= (x+6)(x+6)(x+1)(x+1)
..
The x intercepts
set y=0
(x+6)(x+6)(x+1)(x+1)=0
x=-6,-6,-1,-1
..
The y intercept
set x=0
y=(x+6)(x+6)(x+1)(x+1)=6*6*1*1=36
..
Does the graph cross or touch the x-axis at the smaller(more negative) x intercept?
a- The graph touches x-axis at x=-6
Does the graph cross or touch the x- axis at the larger x intercept?
b- Graph touches x-axis at x=-1